What cycle time would match capacity and demand if demand is 167 units a day, there are 1 shift(s) of 452 minutes each, and workers are given three 45 minute breaks during each shift, one of which is for lunch or dinner? (Round your answer to 2 decimal places.)

Cycle time             minutes per cycle
Answer #1

Demand = 167 Units / Day

1 shift = 452 minutes

There are three breaks of 45 minutes

Net time available for production = 452 - 45 x 3

= 452 - 135

= 317 minutes

Cycle time = Net time available for production / Units to be produced

= 317 / 167

= 1.90 minutes per cycle

Kindly comment if you need further assistance.